# Harmonic analysis of atomic vapor magnetization for in situ calibration of tri-axial time-dependent magnetic field

### v1

@@ III.1 Alignment of the static field along $z$
The coefficients $m_{i}$ represent the ratio between the static field and the rotating one. As the latter must be strong enough to fulfill the adiabatic condition (eq.(1)), $A_{2}$ will be typically small. This can make the method not competitive with the one based on the Larmor frequency minimization.
If the static and the time-dependent field components are generated by different coil sets, misalignments might exist between the $xy$ plane defined by the coils that control the static field and the polarization plane of the time-dependent field. As a consequence, the static field can be made perpendicular either to the $xy$ plane or to the polarization plane. The procedure based on Larmor frequency minimization refers to the former, while the one based on harmonic analysis refers to the latter. The study of coil misalignments goes beyond the scopes of this paper, however we note that the just mentioned feature suggests that using the two methods could help highlight such alignment imperfections.
As shown in the Appendix, non-zero values of $\phi$ and $\varepsilon$ do not contribute to second harmonic terms, making the described procedure robust with respect to imperfect polarization of the rotating field. Indeed (see eqs. (A.3) and (A.4)), the polarization imperfections parametrized by $\phi$ and $\varepsilon$ produce only odd harmonic terms, as discussed in the next.
III.2 Refinement of the field circular polarization on the $xy$ plane
@@
similarly to the case of eq.(4), minimizing $A_{3}$ will lead toward the $\varepsilon=\phi=0$ condition, i.e. to a condition of perfectly circular polarization of the time dependent field.
@@
It is worth examining possible consequences of imperfect perpendicularity of the static field on the accuracy of this procedure. In fact, beside the analyzed second-harmonic terms, a non perpendicular static field causes also third-harmonic signal, mimicking polarization issues. However, as shown in eq.(A.4), $m_{1}$ and $m_{2}$ in a first-order Taylor approximation do not produce such terms, and they only contribute when also the second order (eq.(A.5)) is taken into account. In this sense, the procedure aimed to obtain circular polarization is robust with respect to the $m_{1}=m_{2}=0$ requirement.
IV Validation
In this section we report experimental results validating the proposed methodology. We first investigate the effects of static field misalignments (analyzed in Sec.III.1) and then those of imperfect polarization (analyzed in Sec.III.2).
IV.1 Effects of static field misalignment
@@
We have performed a measurement having aligned the field along $z$ on the basis of Larmor frequency minimization in the absence of any rotating field. The static field inferred from the free precession is about 1 $\mu$T. We have then applied a field of 16 $\mu$T rotating at 3144 Hz, thus the nominal condition is $m_{1}=m_{2}=0,\,m_{3}=1/16$, the relative phase of the oscillating $x,y$ components is set to $\pi/2$ on the basis of the static calibration factor. The recorded signal is reported in Fig.3: no second harmonic peak appear, while a weak third harmonic peak is visible, about 33 dB below the fundamental tone.
@@
Figure 3: Signal generated from atomic magnetization adiabatic following a 16 $\mu$T field rotating at 3.144 kHz on the $xy$ plane, while the static field is the same as for Fig.2, in time (a) and frequency (b) domain, respectively. A second harmonic peak is barely visible, about 45 dB below the fundamental tone.
We have then applied an additional transverse static component that shifts the free precession frequency to 4950 Hz: this corresponds to have $(m_{1},m_{2},m_{3})\approx(0,52,66)\cdot 10^{-3}$, which according to eq.(4) causes a second harmonic peak about 34 dB below the fundamental tone. This expectation is in perfect accordance with the signal analysis shown in Fig.4.
Refer to captionRefer to caption
Figure 4: As in Fig.3, the magnetization adiabatically follows a 16 $\mu$T, 3.144 kHz field rotating on the $xy$ plane. But a transverse component (830 nT in amplitude) is added to the static field along the $z$ direction. Also in this case the signal is plotted in the time (a) and frequency (b) domains. Compared to Fig.3b a clear emergence of a second-harmonic peak occurs. That peak is here about 34 dB below the fundamental tone, in accordance with the prediction of eq.(4).
Noticeably, the measurements shown in Figs.3 and 4 put in evidence also an increase of the decay time with respect to the case of free precession around a static field (Fig.2). Such increase is due to mechanisms emerging when the spin dynamics is driven by a rotating fields much stronger than the static one. These mechanisms are currently under investigation and this phenomenon is not further discussed in this work.
IV.2 Effects of polarization ellipticity
Rotating fields with variable degree of ellipticity are applied to compare theoretical and experimental evaluations of $A_{3}(\varepsilon,\phi)$. In the experiments, the couple $(\varepsilon,\phi)$ is varied around its nominal $(0,0)$ value estimated on the basis of static calibration factors. Discrepancies between static and dynamic calibration factors appear as a translation of the surface $A_{3}(\varepsilon,\phi)$, in particular as a displacement of its minimum.
The plots in Fig.5 shows a 3D representation of calculated $A_{3}(\varepsilon,\phi)$ and corresponding experimental results obtained $m_{1}=m_{2}=0$ (set by minimizing the Larmor frequency in the initial presence of $m_{3}\neq 0$) and with $m_{3}\approx 0$ having minimized the second harmonic peak.
The measurements have been performed with
$\omega=2\pi\cdot 1474$ rad/s, under the application of a rotating field deliberately distorted with the application of variable couples $(\varepsilon,\phi)$ in the range $[-0.15,+0.15]\times[-0.15\mathrm{~{}rad},+0.15\mathrm{~{}rad}]$. The matching is substantial, both in the surface shape and in the absolute values: the minor deviation (about 13%) of the latter is due to a low-pass filtering effect by the transimpedance amplifier of the photodetector.
The same data are then shown as 2D plots in Figs.6a and 6b, in which a displacement of the minimum is barely detectable: $(\varepsilon_{\mathrm{MIN}},\phi_{\mathrm{MIN}})=(8\cdot 10^{-4},0.35\degree)$. Such displacement of the minimum is ascribed to the ellipticity that is obtained when the static calibration factors are used to produce alternating fields.
Higher frequencies of the rotating field cause larger discrepancies from the static calibration, which results in larger displacement of the minimum. The map in Fig.6c is obtained with a 30 $\mu$T field rotating at 4950 Hz. In this case, the minimum of the $A_{3}$ ratio is located in $(\varepsilon_{\mathrm{MIN}},\phi_{\mathrm{MIN}})=(0.0195,-1.7\degree)$
Refer to caption Refer to caption
@@ IV.2 Effects of polarization ellipticity
Figure 6: The theoretical and experimental data reported in Fig.5 are here represented in the 2D maps a) and b), respectively. The map c) is recorded with a 30 $\mu$T field rotating at a higher frequency (4950 Hz). As expected, the higher frequency causes a larger shift of the observed minimum with respect the central point of the map, which corresponds to $(\varepsilon,\phi)=(0,0)$ on the basis of the static calibration. The displacement of the recorded minimum is ($0.358\cdot 10^{-4},$ 0.35°) in case b) and (0.0195, -1.7°) in case c).
@@
The experimental maps confirm the theoretical prediction, and appear with the expected convex surface over wide intervals of the parameters. This suggests that optimization algorithms can be reliably implemented in the experiments to adjust the relative amplitude and phase of the rotating field in order to refine the calibration and to eventually generate a circular polarization at a given frequency.
@@
More generally, the method can be applied to identify the actual phase difference between the drive signal and the magnetic field generated by each coil, i.e. to obtain accurate calibration factors for operating at frequencies where the static calibration are not applicable.
@@ V Conclusion
We have presented an innovative methodology to align static components of a magnetic field and to calibrate its time-dependent components on the basis of in-situ measurements performed with a setup for optical-pumping magnetometry.
The developed procedures are based on the harmonic analysis of polarimetric signals recorded in the presence of an intense, rotating magnetic field and, possibly, of a static one.
We have shown that second harmonic terms can be used to align the static field perpendicularly to the polarization plane of the rotating one. Then we have described a procedure suited to refine the circular polarization of the latter. In typical conditions, the first goal can be better achieved with a traditional approach based on Larmor frequency minimization. In contrast, the method developed to the second end constitutes a very useful tool when relative amplitudes and phases of fast oscillating field components must be precisely assigned.
We have developed and examined the procedure to produce a circularly polarized field on a given plane. Analogous procedures can be implemented on an perpendicular plane and at the diverse frequencies of interest. This would provide a complete set of calibration factors, sufficient to generate three-dimensional fields with arbitrary time-dependence.
@@
A Taylor expansion of $S(t)$ followed by a Fourier analysis leads to determine the signal components at $\omega$ and its multiples. As explained in Sec.III, harmonic terms evidence misalignments of the static field and imperfections of the polarization of the time-dependent field.
It is convenient to normalize the harmonics terms to the fundamental one, to cancel the effects of spurious amplitude fluctuations of $S(t)$ that may occur, e.g., due to fluctuations in the power or in the tuning of the laser sources. Thus we evaluate here in a first order Taylor approximation the terms at $\omega$, $2\omega$ and $3\omega$, which are relevant to the proposed analysis.
@@
with no dependence on $m_{1},m_{2}$: the misalignment terms contribute to the third harmonics only at the second order term, which reads

### v2

@@ III.1 Alignment of the static field along $z$ ($m_{1}=m_{2}=0$ condition)
Similar to the traditional procedure based on Larmor frequency minimization, the transverse components appear quadratically in the quantity $A_{2}$ to be minimized.
If the DC and the AC field components are generated by different coil sets, misalignments might exist between the $xy$ plane defined by the DC coils and the polarization plane of the time-dependent field. As a consequence, the static field can be made perpendicular either to the $xy$ plane or to the polarization plane. The procedure based on Larmor frequency minimization refers to the former, while the one based on harmonic analysis refers to the latter. The study of coil misalignments goes beyond the scope of this paper, however, we note that the just mentioned feature suggests that using the two methods could help highlight imperfect alignment among AC and DC coils.
As shown in the Appendix, non-zero values of $\phi$ and $\varepsilon$ do not contribute to second harmonic terms, making the described procedure robust to imperfect polarization of the rotating field. Indeed (see eqs. (A.3) and (A.4)), the polarization imperfections parametrized by $\phi$ and $\varepsilon$ produce only odd harmonic terms, as discussed in the next.
III.2 Refinement of the field circular polarization on the $xy$ plane ($\phi=\varepsilon=0$ condition)
@@
similarly to the case of eq.(4), minimizing $A_{3}$ will lead toward the $\varepsilon=\phi=0$ condition, i.e. to a condition of perfectly circular polarization of the time-dependent field.
@@
It is worth examining the possible consequences of imperfect perpendicularity of the static field on the accuracy of this procedure. In fact, besides the analyzed second-harmonic terms, a non-perpendicular static field causes also third-harmonic signal, mimicking polarization issues. However, as shown in eq.(A.4), $m_{1}$ and $m_{2}$ in a first-order Taylor approximation do not produce such terms, and they only contribute when also the second order (eq.(A.5)) is taken into account. In this sense, the procedure aimed to obtain circular polarization is robust concerning the $m_{1}=m_{2}=0$ requirement.
IV Experimental validation
IV.1 Static field alignment
@@
We have measured a FPS after having aligned the field along $z$ by Larmor frequency minimization and in the absence of any rotating field. The static field inferred from the FPS is about 1 $\mu$T. We have then applied a field of 16 $\mu$T rotating at 3144 Hz, thus the nominal condition is $m_{1}=m_{2}=0,\,m_{3}=1/16$, the amplitudes and the relative phase of the oscillating $x,y$ components are set to $\pi/2$ based on the static calibration factors. The recorded signal is reported in Fig.4: just a weak second harmonic peak appears, and third harmonic peak is visible, about 33 dB below the fundamental tone.
Figure 4: Signal generated from atomic magnetization adiabatic following a 16 $\mu$T field rotating at 3.144 kHz on the $xy$ plane, while the static field is the same as for Fig.3, in time (a) and frequency (b) domain, respectively. A second harmonic peak is barely recognizable, about 50 dB below the fundamental tone. The red, blue, orange, and green lines indicate the field rotation frequency and its $2^{\mathrm{nd}}$, $3^{\mathrm{rd}}$, and $4^{\mathrm{th}}$ harmonics, respectively.
We have then applied an additional transverse static component that shifts the free precession frequency to 4950 Hz: this corresponds to have $(m_{1},m_{2},m_{3})\approx(0,52,66)\cdot 10^{-3}$, which according to eq.(4) causes a second-harmonic peak about 34 dB below the fundamental tone. This expectation is in perfect accordance with the spectrum shown in Fig.5b. For comparison, the condition of Fig.4, with the second-harmonic peak 50 dB weaker than the fundamental tone (hence $A_{2}\approx 10^{-5}$) gives $m_{1}^{2}+m_{2}^{2}\approx 8\cdot A_{2}\approx 8\cdot 10^{-5}$. Despite this small value, as $m_{3}$ is also small, this corresponds to a $8\degree$ misalignment. This value would decrease down to $2.6\degree$ in case of a 60 dB ratio, i.e. if the second-harmonic peak were at the level of the noise floor.
Figure 5: As in Fig.4, the magnetization adiabatically follows a 16 $\mu$T, 3.144 kHz field rotating on the $xy$ plane. But a transverse component (830 nT in amplitude) is added to the static field along the $z$ direction. Also in this case the signal is plotted in the time (a) and frequency (b) domains. Compared to Fig.4b a second-harmonic peak emerges, in quantitative accordance with the prediction of eq.(4). The red, blue, orange, and green lines indicate the field rotation frequency and its $2^{\mathrm{nd}}$, $3^{\mathrm{rd}}$, and $4^{\mathrm{th}}$ harmonics, respectively.
Noticeably, the measurements shown in Figs.4 and 5 put in evidence also an increase of the decay time compared to the case of free precession around a static field (Fig.3). Such increase is due to mechanisms emerging when the spin dynamics is driven by a rotating fields much stronger than the static one. These mechanisms are currently under investigation and this phenomenon is not further discussed in this work.
IV.2 Effects of polarization ellipticity
Rotating fields with variable degrees of ellipticity are applied to compare theoretical and experimental evaluations of $A_{3}(\varepsilon,\phi)$. In the experiments, the couple $(\varepsilon,\phi)$ is varied around its nominal $(0,0)$ value estimated on the basis of static calibration factors. Discrepancies between static and dynamic calibration factors appear as a translation of the surface $A_{3}(\varepsilon,\phi)$, in particular as a displacement of its minimum.
The plots in Fig.6 shows a 3D representation of calculated $A_{3}(\varepsilon,\phi)$ and corresponding experimental results obtained with $m_{1}=m_{2}=0$.
The measurements have been performed with $\omega=2\pi\cdot 1474$ rad/s, under the application of a rotating field deliberately distorted with the application of variable couples $(\varepsilon,\phi)$ in the range $[-0.15,+0.15]\times[-0.15\mathrm{~{}rad},+0.15\mathrm{~{}rad}]$. The matching is substantial, both in the surface shape and in the absolute values: the minor deviation (about 13%) of the latter is due to a low-pass filtering effect by the transimpedance amplifier of the photodetector.
The same data are then shown as 2D plots in Figs.7a and 7b, to facilitate the visualization of the displacement of the minimum. In this case, such displacement is barely detectable: $(\varepsilon_{\mathrm{MIN}},\phi_{\mathrm{MIN}})=(8\cdot 10^{-4},0.35\degree)$.
Higher frequencies of the rotating field cause larger discrepancies from the static calibration and result in a larger ellipticity and consequently larger displacements of the minimum. The map in Fig.7c is obtained with a 30 $\mu$T field rotating at 4950 Hz. In this case, the minimum of the $A_{3}$ ratio is located in $(\varepsilon_{\mathrm{MIN}},\phi_{\mathrm{MIN}})=(0.0195,-1.7\degree)$, and it correspondingly decreases by about 15 dB, down to the level set by the noise floor.
@@ IV.2 Effects of polarization ellipticity
Figure 7: The theoretical and experimental data reported in Fig.6 are here represented in the 2D maps a) and b), respectively. The experimental map c) is recorded with a 30 $\mu$T field rotating at a higher frequency (4950 Hz). As expected, the higher frequency causes a larger displacement of the observed minimum from the map’s central point, which corresponds to the nominal $(\varepsilon,\phi)=(0,0)$ based on the static calibration. The displacement of the recorded minimum is ($0.358\cdot 10^{-4},$ 0.35°) in case b) and (0.0195, -1.7°) in case c).
@@
The experimental maps confirm the theoretical prediction and appear with the expected convex surface over wide intervals of the parameters. This suggests that optimization algorithms could be reliably implemented to adjust the relative amplitude and phase of the rotating field to refine the coil calibration and eventually generate a circular polarization at any assigned frequency.
@@
More generally, the method can be applied to identify the phase difference between the drive signal and the magnetic field generated by each coil making it possible to generate fields with an arbitrary configuration.
@@ V Conclusion
We have introduced a methodology to align static components of a magnetic field and to calibrate its time-dependent components based on in-situ measurements performed with a setup for optical-pumping magnetometry. The procedures are based on the harmonic analysis of polarimetric signals recorded in the presence of an intense, rotating magnetic field and, possibly, of a static one.
We have shown that second harmonic terms can be used to point out the presence of DC field components lying on the polarization plane of the rotating one. Then we have described a procedure suited to refine the circular polarization of the latter. In typical conditions, the first goal can also be pursued with a traditional approach based on Larmor frequency minimization. In contrast, the method developed for polarization refinement constitutes a useful tool to retrieve accurate calibration factors, which enable a precise set of relative amplitudes and phases of fast-oscillating field components.
We have developed and examined the procedure to produce a circularly polarized field on a given plane. Analogous procedures can be implemented on a perpendicular plane and at the diverse frequencies of interest. This would provide a complete set of calibration factors sufficient to generate three-dimensional fields with arbitrary configuration and time dependence.
@@
A multivariate Taylor expansion of $S(t)$ in the small variables $m_{1},m_{2},\epsilon$ and $\phi$ followed by a Fourier analysis leads to determining the signal components at $\omega$ and its multiples.
It is convenient to normalize the harmonics terms to the fundamental one, to cancel the effects of spurious amplitude fluctuations of $S(t)$ that may occur, e.g., due to power or tuning fluctuations of the laser sources. Thus we evaluate here in a first-order Taylor approximation the terms at $\omega$, $2\omega$ and $3\omega$, which are relevant to the proposed analysis.
@@
with no dependence on $m_{1},m_{2}$: the misalignment terms contribute to the third harmonics only at the second order of the Taylor expansion, which reads